OpenStudy (anonymous):

Really need help with this! The social studies teacher is giving a test worth 50 points, and containing 20 questions. There are 2 point and 4 point questions on the test. How many of each type of question are on the test? Show your work.

OpenStudy (imstuck):

you need a system of equations here, one equation that relates the number of each type of question to the total number of questions, which is 20; you then need another equation that represents how much each type of question is worth to the total number of points which is 50. Like this:

OpenStudy (anonymous):

2x+4y=50 x+y=20

OpenStudy (imstuck):

The number of questions is represented by x + y = 20 (x being the number of 2 pt questions and y being the number of 4 pt questions). The amount of points is represented by 2x + 4y = 50 (because each x question is worth 2 points and each y question is worth 4 points.) Solve these simultaneously.

OpenStudy (anonymous):

so it would be x=15 and y=5?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks!